Help me please, thank you.
Graph the quadratic
y+8+3x^2=12x
show all work if some thanks !

Answer by ewatrrr(24785)


Hi
y+8+3x^2=12x
y = -3x^2 + 12x - 8 P(0,-8) and (1,1) on this parabola opening downward -3 < 0
Completing the Square
y = -3(x-2)^2 + 12 - 8
y = -3(x-2)^2 + 4 V(2,4) line of symmetry x = 2